the sequence shown below is defined using a recursion formula. write the first four terms of the sequence.
a1 13 and an=an-1+8 for n>=2

Answer by amarjeeth123(569)

a1=13
an=an-1+8
a2=a1+8=21
a3=a2+8=29
a4=a3+8=37